All techniques -Decision among mutually exclusive investments Pound Industries is attempting to select the best of three mutually exclusive projects. The initial investment and after-tax cash inflows associated with these projects are shown in the following table. Cash flows Initial investment (CF) Cash inflows (CF), t-1 to 5 $100,000 $30,000 Project A Project B $120,000 $41,000 Poject C $130,000 $42,500 a. Calculate the payback period for each project. b. Calculate the net present value (NPI) of each project, assuming that the firm has a cost of capital equal to 13%. c. Calculate the internal rate of return (IRR) for each project. d. Indicate which project you would recommend.
Answer #1

a. Payback period for each project= Cash Outlay/ Annual Cash Inflow

Project A =100000/30000= 3.33 years

Project B = 120000/41000= 2.92 years

Project C= 130000/42500= 3.05 years

b. NPV = Discounted Cash Flow -Initial Investment

Year Cash Flow Discounting Factor Present Value
0 100000 0 (100000)
1-5 30000 3.5172 105516
NPV 5516

Project A- NPV = 1,05,516 *-1,00,000 =5516   

Year Cash Flow Discounting Factor Present Value
0 120000 0 (120000)
1-5 41000 3.5172 144206
NPV 24206

Project B-NPV = 1,44,206-1,20,000= 24206

Year Cash Flow Discounting Factor Present Value
0 130000 0 (130000)
1-5 42500 3.5172 149482
NPV 19482

Project C- NPV = 149482- 1,30,000= 19482

(c) IRR of the project

Internal rate of return is a discount rate that makes the net present value (NPV) of all cash flows from a particular project equal to zero.

Project A

Year Cash Flow Discount factor at 10% PV Discount factor at 17% PV
0 100000 1 (100000) 1 (100000)
1-5 30000 3.791 113723 4.273

128205

NPV at 10% 13723

NPV at 17% 28205

IRR= L + NPV (L) * (H-L)

NPV (L)-NPV (H)

IRR= 10+ 13723 *7 = 10.36%

13723- 28205

Project B

Year Cash Flow Discount factor at 10% PV Discount factor at 17% PV
0 120000 1 (120000) 1 (120000)
1-5 41000 3.791 155422 4.273

175213

NPV at 10% =35422

NPV at 17% =55213

IRR= L + NPV (L) * (H-L)

NPV (L)-NPV (H)

IRR= 10+ 35422 *7 = 22.52%

55213- 35422

Project B

Year Cash Flow Discount factor at 10% PV Discount factor at 17% PV
0 130000 1 (130000) 1 (130000)
1-5 42500 3.791 161108 4.273 181623
31108 51623

NPV at 10% =31108

NPV at 17% =51623

IRR= L + NPV (L) * (H-L)

NPV (L)-NPV (H)

IRR= 10+ 31108 *7 = 20.61 %

20515

Project B is recommendable as it is having lower pay back period, highest NPV and highest IRR among all 3 projects.